What do you understand by electrostatic potential energy? Calculate the
electrostatic potential energy for the system of charges shown below. Take q = 4μC
and a = 2 cm.
1
Expert's answer
2017-06-29T10:53:07-0400
The electrostatic potential energy is is the energy a charge has due to its position relative to other charges.
E_pot=(kq^2)/a=((9∙〖10〗^9 ) (4∙〖10〗^(-6) )^2)/((0.02))=7.2 J.